OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (6, 2) and has a slope of 3? y = x +

pooja195 (pooja195):

you need to use slop intercept form to find your b value so y-y1=m(x-x1)

OpenStudy (cwrw238):

here x1 = 6, y1 = 2 and m = 3 just plug these in and simplify

pooja195 (pooja195):

y-2=3(x-6) can you solve that?

OpenStudy (anonymous):

y=3x-16

pooja195 (pooja195):

yep thats correct!! good job WELCOME TO OPENSTUDY! :) Don't hesitate to ask questions. :)

OpenStudy (anonymous):

thxs